Solve each system using substitution 2x = y-4 and -3y = x - 5

Answer by tommyt3rd(5050)

2x+4 = y
and so
-3(2x+4) = x - 5
-6x-12=x-5
12-5=x+6x

x=-7/7=-1
y=2